OpenStudy (anonymous):

Find the following product:(x-2)(x+2) A. x^2+4x-4 B.x^2-4 C.x^2+4 D.x^2-4x-4

OpenStudy (anonymous):

Use FOIL to expand. \[x ^{2}+2x-2x-2^{2}\]

OpenStudy (anonymous):

b

OpenStudy (anonymous):

(x-2)(x+2) Foil: \[x\ \times\ x\ =x ^{2}\] \[x\ \times\ 2 = 2x\] \[-2\ \times\ x = -2x\] \[-2\ \times\ 2 = -4\] Answer: \[x ^{2}+2x-2x-4\]Final Answer\[x ^{2}-4\] Since 2x-2x cancel each other out. So the answer is b.